# Supertranslation ambiguity in post-Minkowskian expansion

### v2

@@
Abstract: The supertranslation ambiguity of angular momentum is a long-standing problem in general relativity. In the context of post-Minkowskian expansion, the supertranslation ambiguity arises the puzzle of the angular momentum loss to start at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ in gravitational scattering. In this paper, we propose a generic prescription to fix the supertranslation ambiguity at the linear order in post-Minkowskian expansion which will uniquely determine the angular momentum loss. The proposal is self-contained in the post-Minkowskian expansion and involves only null infinity data.
@@ 1 Introduction
The success of direct observation of gravitational waves opens new avenues for gravitational wave (GW) astronomy [1]. The accurate theoretical modeling of gravitational scattering is crucial for the GW detectors [2]. Currently, the waveform of gravitational radiation is mainly derived in expansion based on successive approximations in powers of the Newton’s constant $G$, namely the post-Minkowskian (PM) expansion. While, the linear nature of the PM expansion used to raise huge disputes that if the gravitational radiation was just an artifact of linearization up until the 1960s. The dispute is settled by Bondi and collaborators by formulating Einstein equation as a characteristic initial value problem at null infinity [3, 4]. In Bondi’s framework, the gravitational radiation is characterized by the news functions and the mass of the system always decreases whenever news functions exist, which eventually confirms the existence of gravitational waves in the full Einstein theory.
@@
A surprising result from Bondi’s framework [3, 4] is that the asymptotic symmetry group at null infinity is not the Poincaré group, but the semidirect sum of supertranslations, an infinite dimensional Abelian group which extends the translations, with the global conformal group of the celestial sphere that is isomorphic to the Lorentz group. Though, the enhancement from translation to supertranslation has several remarkable applications recently, see, e.g., the comprehensive review in [5], it also brings puzzles in general relativity. A long-standing problem is the supertranslation ambiguity in defining angular momentum [6, 7]. In the context of the PM expansion, the supertranslation ambiguity causes the puzzle of angular momentum loss in gravitational scattering [8]. The leading order of the angular momentum loss can be at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ depending on different gauge choices under the supertranslation [9, 10, 11, 12, 13, 14, 15, 8, 16, 17, 18, 19, 20, 21, 22]. Angular momentum loss formula at null infinity and its relation to the change of angular momentum of the source are certainly of utmost importance for higher precision PM computations, which lead to the recent investigations from various perspectives, including the binary systems [23, 24], supertranslation invariant definition of angular momentum at null infinity [25, 26, 27, 28, 29, 30, 31] and at spatial infinity [32, 33], supertranslation frame-fixing at the far past of the future null infinity [8, 34], supertranslation frame-fixing from covariance [35], memory effect [36, 37], and flux balance law [37, 38].
In this paper, we propose a self-contained prescription in the PM expansion to fix the supertranslation ambiguity at the linear order which can uniquely determine the angular momentum loss. The resolution is encoded in the linear nature of the PM expansion and involves only null infinity data. In Einstein theory, the supertranslation freedom is intrinsic in the sense that it is associated with the propagating degree of freedom. One can not fix the gauge choice associated to the supertranslation freedom. Because gravitational radiation will change the gauge choice in any case [7], see also [39] for a modern interpretation. When the Einstein equation is expanded about the Newton’s constant $G$, one can recover a linearized theory at each order with its own gauge structure which is inherited from the diffeomorphism invariance of the Einstein theory. When one also expands a generic transformation of coordinates system in $G$, the transformation at each order of $G$ is the gauge transformation of the linearized theory at the corresponding order. Adapting the Bondi’s framework to the PM expansion, one can recover a supertranslation at each order of $G$. A fine tuning structure arises during the linearization. The propagating degree of freedom is at ${\cal O}(G^{2})$. While the disagreement of angular momentum loss is relevant to the first order supertranslation. This is a conceptually crucial feature since we can fix the first order supertranslation to resolve the angular momentum loss puzzle without touching the propagating degree of freedom at ${\cal O}(G^{2})$.
Consider that the spacetime is sourced by a system of pointlike bodies. The relativistic definitions of 4-momentum and angular momentum from the stress tensor are the same as the ADM definitions of the linear order metric [8]. Though, the linear metric in the Bondi gauge has first order supertranslation dependence [8, 40], the 4-momentum and angular momentum defined at null infinity are independent of the first order supertranslation and coincide with the relativistic definitions [40]. It is important to point out that the zeroth order supertranslation was not involved at all in [8, 40]. One can think of that the coincidence of the charges at null infinity and the relativistic definition of charges in the bulk rules out the zeroth order supertranslation. Then the key idea we will implement is to use the charges at the second order to control the first order supertranslation. Naively, this is impossible since there is no conservation law at the second order due to the gravitational radiation. Nevertheless, we can consider part of the second order solution. The source of the linearized equations of motion at the second order consists of two parts. The corrected stress tensor of the pointlike bodies due to the linear order metric and the effective stress tensor from the first order metric. Since the theory is linear, we can study those two parts independently. If one ignores the modification of the motion of the pointlike bodies and considers only the effective stress tensor, which we would refer to as the self-part, [^1] The self-part in our terminology includes the self-terms and cross-terms in a two pointlike objects system investigated in [41]. the relativistic definition of 4-momentum and angular momentum will not be corrected from the next-to-leading order in the expansion of $G$. Hence, it is reasonable to require that the second order metric of the self-part at null infinity should have vanishing 4-momentum and angular momentum. This is the criterion we propose to fix the supertranslation ambiguity at the first order.
Since the waveform at the second order from the source of arbitrary numbers of pointlike bodies is not yet known, nor the self-part of this system, we can not test our prescription in the generic case. We will use an alternative approach to prove that the first order supertranslation freedom will be ruled out by our prescription. We apply the asymptotic analysis to derive the generic second order solution with the effective stress tensor in series expansion near null infinity. Then the criteria of vanishing of 4-momentum and angular momentum will fix the initial data of the second order asymptotic solution. We then compute the transformation laws of the 4-momentum and angular momentum under the first order supertranslation. We show that a generic first order supertranslation will change the next-to-leading order angular momentum. The vanishing conditions of 4-momentum and angular momentum finally fix the first order supertranslation.
2 Self-part at the second order
We linearize Einstein theory around the background Minkowski spacetime to the second order $g_{\mu\nu}=\eta_{\mu\nu}+GH_{\mu\nu}+G^{2}h_{\mu\nu}$. The inverse metric up to the second order is
$g^{\mu\nu}=\eta^{\mu\nu}-GH^{\mu\nu}+G^{2}(H^{\mu}_{\alpha}H^{\nu\alpha}-h^{\mu\nu}),$ (1)
where the indices are now raised by $\eta^{\mu\nu}$. The equations of motion for the self-part at the second order are [42, 41] (see also Appendix A)
$E_{\mu\nu}\equiv\left[\bar{\nabla}_{\tau}\bar{\nabla}_{\mu}h^{\tau}_{\nu}+\bar{\nabla}_{\tau}\bar{\nabla}_{\nu}h^{\tau}_{\mu}-\bar{\nabla}^{2}h_{\mu\nu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h\right]\\ -\eta_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}h^{\alpha\beta}-\bar{\nabla}^{2}h\right)=T_{\mu\nu}^{E},$ (2)
where $h=\eta^{\mu\nu}h_{\mu\nu}$, $H=\eta^{\mu\nu}H_{\mu\nu}$, and the effective stress tensor is given by
$T^{E}_{\mu\nu}=\eta_{\mu\nu}\bigg{[}H^{\alpha\beta}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H+\bar{\nabla}^{2}H_{\alpha\beta}-2\bar{\nabla}^{\tau}\bar{\nabla}_{\alpha}H_{\tau\beta}\right)-\frac{1}{2}\bar{\nabla}_{\beta}H_{\alpha\tau}\bar{\nabla}^{\tau}H^{\alpha\beta}\\ +\bar{\nabla}_{\beta}H\left(\bar{\nabla}_{\alpha}H^{\alpha\beta}-\frac{1}{4}\bar{\nabla}^{\beta}H\right)-\bar{\nabla}_{\alpha}H^{\alpha\beta}\bar{\nabla}^{\tau}H_{\tau\beta}+\frac{3}{4}\bar{\nabla}_{\alpha}H^{\tau\beta}\bar{\nabla}^{\alpha}H_{\tau^{\beta}}\bigg{]}\\ +H^{\tau\alpha}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\alpha}\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\alpha}\bar{\nabla}_{\tau}H_{\mu\nu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}H_{\tau\alpha}\right)\\ +H_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H^{\alpha\beta}-\bar{\nabla}^{2}H\right)+\bar{\nabla}_{\alpha}H^{\alpha\tau}\left(\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\tau}H_{\mu\nu}\right)\\ -\frac{1}{2}\bar{\nabla}_{\mu}H_{\tau\alpha}\bar{\nabla}_{\nu}H^{\tau\alpha}+\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\tau}H_{\mu}^{\alpha}-\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\alpha}H_{\mu}^{\tau}\\ -\frac{1}{2}\bar{\nabla}^{\alpha}H\left(\bar{\nabla}_{\mu}H_{\nu\alpha}+\bar{\nabla}_{\nu}H_{\mu\alpha}-\bar{\nabla}_{\alpha}H_{\mu\nu}\right).$ (3)
We use $\bar{\nabla}$ to denote the derivative in the Minkowski spacetime. The full first order metric $H_{\mu\nu}$ in Bondi gauge up to relevant orders in asymptotic expansions is obtained in [40],
$H_{uu}=$ $\frac{2GM_{b}}{r}+{\cal O}(\frac{1}{r^{2}}),\quad\quad M_{b}=-\sum{\frac{m}{(v_{\mu}n^{\mu})^{3}}},$ (4)
$H_{ur}=$ ${\cal O}(\frac{1}{r^{2}}),$ (5)
$H_{AB}=$ $rGC_{AB}+{\cal O}(1),\quad\quad C_{AB}=\left[2D_{{A}}D_{{B}}-{\gamma}_{{A}{B}}D_{{C}}D^{{C}}\right]\left[\bar{\beta}(x^{A})-f(x^{A})\right],$ (6)
$f(x^{A})=2\sum m(n\cdot v)\log(-(n\cdot v)),$
$H_{uA}=$ $\frac{G}{2}D^{B}C_{AB}+\frac{2}{3r}\left[uD_{A}\left(GM_{b}\right)+GN_{A}\right]+{\cal O}(\frac{1}{r^{2}}),$ (7)
$N_{A}=D_{A}\left\{\frac{3}{2}\sum{m\left[-\frac{(n^{\mu}c_{\mu})^{2}}{(n^{\nu}v_{\nu})^{3}}+\frac{v^{\mu}c_{\mu}n^{\nu}c_{\nu}}{(n^{\sigma}v_{\sigma})^{2}}+\frac{v^{\mu}c_{\mu}}{(n^{\nu}v_{\nu})^{2}}+\frac{1}{3}\frac{n^{\mu}c_{\mu}}{(n^{\nu}v_{\nu})^{3}}\right]}\right\}$
$\quad\quad-\sum{m\frac{2}{(n^{\nu}v_{\nu})^{3}}D_{A}(n^{\mu}c_{\mu})},$

### v3

@@
Abstract: The supertranslation ambiguity of angular momentum is a long-standing problem in general relativity, which arises the puzzle of the angular momentum loss in the post-Minkowskian (PM) expansion. The angular momentum loss in gravitational scattering is of order $G^{2}$ in the linear response theory but of order $G^{3}$ in the amplitudes-based methods. In this paper, we propose a generic prescription to fix the supertranslation ambiguity in the PM expansion which will uniquely determine the angular momentum loss. The proposal is self-contained and involves only the null infinity data.
@@ 1 Introduction
The success of direct observation of gravitational waves opens new avenues for gravitational wave astronomy [1]. The accurate theoretical modeling of gravitational scattering is crucial for the gravitational wave detectors [2]. Currently, the waveform of gravitational radiation is mainly derived in expansion based on successive approximations in powers of some small parameters, such as the inverse of the speed of light $1/c$ or the Newton’s constant $G$, namely the post-Newtonian or post-Minkowskian expansions, respectively, see, e.g., the reviews in [3, 4]. A common feature of those approximate methods is that the Einstein equation is represented by a linear wave equation at each order in the expansion on the Minkowski or curved background and the non-linear effect of the Einstein theory is encoded in the effective stress tensor term, which raises huge disputes that if the gravitational radiation was just an artifact of the linearization from the expansion up until the 1960s. The dispute is settled by Bondi and collaborators by formulating the Einstein equation as a characteristic initial value problem at null infinity [5, 6]. In the Bondi’s framework, the gravitational radiation is characterized by the news functions in the asymptotic metric and the mass of the system always decreases whenever news functions exist, which eventually confirms the existence of gravitational waves in the full Einstein theory.
@@
A surprising result from the Bondi’s framework [5, 6] is that the asymptotic symmetry group at the null infinity is not the Poincaré group, but the semidirect sum of supertranslations, an infinite dimensional Abelian group which extends the translations, with the global conformal group of the celestial sphere that is isomorphic to the Lorentz group. The enhancement from translation to supertranslation has several remarkable applications recently, see, e.g., the comprehensive review in [7], but it also brings puzzles in general relativity. A long-standing problem is the supertranslation ambiguity in defining angular momentum [8, 9]. In the context of the PM expansion, the supertranslation ambiguity causes the puzzle of angular momentum loss in gravitational scattering [10]. The lowest order of the angular momentum loss can be at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ depending on different gauge choices under the supertranslation [11, 12, 13, 14, 15, 16, 17, 10, 18, 19, 20, 21, 22, 23, 24]. The angular momentum loss formula at null infinity and its relation to the change of angular momentum of the source are certainly of utmost importance for higher precision PM computations, which leads to the intensive investigations recently from various perspectives, including the detailed analysis of the compact binary coalescence systems [25, 26], supertranslation invariant definitions of the angular momentum at null infinity [27, 28, 29, 30, 31, 32, 33] and at spatial infinity [34, 35], supertranslation frame-fixing from the matching condition [10, 36] and from covariance [37], memory effect [38, 39], and flux balance law [39, 40].
In this paper, we propose to use asymptotic charges in the PM expansion to fix the supertranslation ambiguity. Our proposal is self-contained at the null infinity and does not involve any data from the spatial infinity. The resolution is encoded in the linear nature of the wave equation in the PM expansion. When the Einstein equation is expanded about the Newton’s constant $G$, one can recover a linear wave equation at each order of $G$ which is sourced by the stress tensor from the matter fields at each order and the effective stress tensor consisting of the perturbative metric in the lower orders of $G$ that reflects the non-linear feature of the Einstein theory. At each order of $G$, the sourced wave equation is invariant under gauge transformations which are inherited from the diffeomorphism invariance of the Einstein theory. In particular, there is a supertranslation at each order of $G$ when combining the PM expansion with the Bondi’s framework [41].
Consider that the full metric is expanded around the background Minkowski spacetime to the second order in $G$ as $g_{\mu\nu}=\eta_{\mu\nu}+GH_{\mu\nu}+G^{2}h_{\mu\nu}$. Correspondingly, a diffeomorphism transformation in the infinitesimal version can be also expanded as
$\delta g_{\mu\nu}\equiv{\cal L}_{\xi}g_{\mu\nu}={\cal L}_{\xi}(\eta_{\mu\nu}+GH_{\mu\nu}+G^{2}h_{\mu\nu}).$ (1)
If one also expands the symmetry parameter in $G$ as $\xi=G\xi^{(1)}+G^{2}\xi^{(2)}$, [^1] One should consider $G$ as a dimensionless parameter. Otherwise, it should be $GM_{c}/(c^{2}r_{c})$, where $M_{c}$ is a mass parameter and $r_{c}$ is a radial parameter. Here, we introduce the $G$ expansion in the symmetry transformation as an effective way of deriving gauge transformation of the theory at each order of $G$, see also [10] for a similar treatment. the induced infinitesimal gauge transformation at order $G$ is
$\delta H_{\mu\nu}={\cal L}_{\xi^{(1)}}\eta_{\mu\nu},$ (2)
which is a pure gauge transformation that will not change the curvature tensor at order $G$. Correspondingly, the supertranslation at order $G$ will not change the leading order of the linear and angular momentum in the expansion of $G$ [42]. [^2] The supertranslation at order $G$ will affect the angular momentum loss at the leading order [10]. But the leading order angular momentum, which is different from the leading order angular momentum loss, is independent of the order $G$ supertranslation. At order $G^{2}$, the equations of motion are invariant under the transformation
The first piece on the right hand side is the standard gauge transformation. The second piece is relevant to the gauge transformation at order $G$, which is not necessary to be a pure gauge transformation at order $G^{2}$. The perturbative metric at order $G^{2}$ satisfies the Fierz-Pauli equation sourced by two parts, the stress tensor of the matter fields $T_{\mu\nu}^{M}$ with corrections from the perturbative metric at order $G$ and the effective stress tensor $T_{\mu\nu}^{E}$ constructed purely from the order $G$ metric,
$E_{\mu\nu}\equiv$ [¯$\left[\bar{\nabla}_{\tau}\bar{\nabla}_{\mu}h^{\tau}_{\nu}+\bar{\nabla}_{\tau}\bar{\nabla}_{\nu}h^{\tau}_{\mu}-\bar{\nabla}^{2}h_{\mu\nu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h\right]-\eta_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}h^{\alpha\beta}-\bar{\nabla}^{2}h\right)$
$=$ $T_{\mu\nu}^{E}+T_{\mu\nu}^{M},$ (4)
where $h=\eta^{\mu\nu}h_{\mu\nu}$, $H=\eta^{\mu\nu}H_{\mu\nu}$, and we use $\bar{\nabla}$ to denote the derivative in the Minkowski spacetime. The Fierz-Pauli equation is linear, which guarantees that the metric can be a superimposition as $h_{\mu\nu}=h_{\mu\nu}^{S}+h_{\mu\nu}^{M}$ where the first piece $h_{\mu\nu}^{S}$, which we would refer to as the self-part, [^3] The self-part in our terminology includes the self-terms and cross-terms in a two pointlike objects system investigated in [43]. The split of the metric can be obviously performed in the Cartesian coordinates system. The equations of motion for both $h^{S}_{\mu\nu}$ and $h^{M}_{\mu\nu}$ are in a covariant way. Hence, one can write both equations in any other coordinates system. satisfies the wave equation sourced by $T_{\mu\nu}^{E}$ and the second piece $h_{\mu\nu}^{M}$ satisfies the wave equation sourced by $T_{\mu\nu}^{M}$. One can verify that both $T_{\mu\nu}^{E}$ and $T_{\mu\nu}^{M}$ are conserved. Hence, the equations for both $h_{\mu\nu}^{S}$ and $h_{\mu\nu}^{M}$ are invariant under the transformation represented by the first piece on the right hand of (3). However, the second piece on the right hand of (3) is not an independent gauge transformation for both $h_{\mu\nu}^{S}$ and $h_{\mu\nu}^{M}$.
Considering only the self-part, we will show that the next-to-leading order angular momentum in the expansion of $G$ is mainly controlled by the order $G$ supertranslation. Thus, the prescription we will propose to fix the order $G$ supertranslation, hence to resolve the supertranslation ambiguity of the angular momentum loss in the PM expansion, is from a physical requirement of the next-to-leading order angular momentum from the self-part. In Einstein theory, the supertranslation freedom is intrinsic in the sense that it is associated with the propagating degree of freedom. In other words, fixing the gauge choice associated to the supertranslation freedom is equivalent to turning off the gravitational radiation [9], see also [44] for a modern interpretation. It is well known that the propagating degree of freedom represented by the news functions starts at order $G^{2}$ in the PM expansion, see, e.g., [10, 11]. Thus we can fix the supertranslation at order $G$ without touching the propagating degree of freedom at order $G^{2}$, which is a fine tuning structure arising in the PM expansion. The key point now is what the physical requirement of the next-to-leading order angular momentum for the self-part should be.
2 The resolution to supertranslation ambiguity
The criteria we propose to fix the order $G$ supertranslation is that the self-part $h_{\mu\nu}^{S}$ does not contribute to the next-to-leading order linear and angular momentum. The effective stress tensor at order $G^{2}$ is given by [43, 45]
$T^{E}_{\mu\nu}=$ $\eta_{\mu\nu}\bigg{[}H^{\alpha\beta}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H+\bar{\nabla}^{2}H_{\alpha\beta}-2\bar{\nabla}^{\tau}\bar{\nabla}_{\alpha}H_{\tau\beta}\right)-\frac{1}{2}\bar{\nabla}_{\beta}H_{\alpha\tau}\bar{\nabla}^{\tau}H^{\alpha\beta}$
+¯$+\bar{\nabla}_{\beta}H\left(\bar{\nabla}_{\alpha}H^{\alpha\beta}-\frac{1}{4}\bar{\nabla}^{\beta}H\right)-\bar{\nabla}_{\alpha}H^{\alpha\beta}\bar{\nabla}^{\tau}H_{\tau\beta}+\frac{3}{4}\bar{\nabla}_{\alpha}H^{\tau\beta}\bar{\nabla}^{\alpha}H_{\tau^{\beta}}\bigg{]}$
+¯$+\bar{\nabla}_{\alpha}H^{\alpha\tau}\left(\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\tau}H_{\mu\nu}\right)$
$+H^{\tau\alpha}\big{(}\bar{\nabla}_{\alpha}\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\alpha}\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\alpha}\bar{\nabla}_{\tau}H_{\mu\nu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}H_{\tau\alpha}\big{)}$
$+H_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H^{\alpha\beta}-\bar{\nabla}^{2}H\right)-\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\alpha}H_{\mu}^{\tau}-\frac{1}{2}\bar{\nabla}_{\mu}H_{\tau\alpha}\bar{\nabla}_{\nu}H^{\tau\alpha}$
+¯$+\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\tau}H_{\mu}^{\alpha}-\frac{1}{2}\bar{\nabla}^{\alpha}H\left(\bar{\nabla}_{\mu}H_{\nu\alpha}+\bar{\nabla}_{\nu}H_{\mu\alpha}-\bar{\nabla}_{\alpha}H_{\mu\nu}\right).$ (5)
The order $G$ metric $H_{\mu\nu}$ of the spacetime sourced by a system of pointlike bodies is known and transformed into the Bondi’s framework recently [42]. Up to relevant orders in $1/r$ series expansion near null infinity, the metric is given by
$H_{uu}=\frac{2M_{b}}{r}+{\cal O}(\frac{1}{r^{2}}),\quad\quad M_{b}=-\sum{\frac{m}{(v_{\mu}n^{\mu})^{3}}},$ (6)
$H_{ur}={\cal O}(\frac{1}{r^{2}}),\quad\quad H_{AB}=rC_{AB}+{\cal O}(1),$ (7)
$C_{AB}=\left[2D_{{A}}D_{{B}}-{\gamma}_{{A}{B}}D_{{C}}D^{{C}}\right]\left[\bar{\beta}(x^{A})-f(x^{A})\right],$ (8)
$\quad\quad f(x^{A})=2\sum m(n\cdot v)\log(-(n\cdot v)),$
$H_{uA}=\frac{1}{2}D^{B}C_{AB}+\frac{2}{3r}\left[uD_{A}\left(M_{b}\right)+N_{A}\right]+{\cal O}(\frac{1}{r^{2}}),$ (9)
$N_{A}=D_{A}\bigg{\{}\frac{3}{2}\sum m\bigg{[}-\frac{(n^{\mu}c_{\mu})^{2}}{(n^{\nu}v_{\nu})^{3}}+\frac{v^{\mu}c_{\mu}n^{\nu}c_{\nu}}{(n^{\sigma}v_{\sigma})^{2}}+\frac{v^{\mu}c_{\mu}}{(n^{\nu}v_{\nu})^{2}}$
$\quad\quad+\frac{1}{3}\frac{n^{\mu}c_{\mu}}{(n^{\nu}v_{\nu})^{3}}\bigg{]}\bigg{\}}-\sum{m\frac{2}{(n^{\nu}v_{\nu})^{3}}D_{A}(n^{\mu}c_{\mu})},$ (10)